# A Universal Framework for Horizon-Scale Tests of Gravity with Black Hole Shadows

### v1

@@ Abstract
Remarks.—In summary, we have constructed a general and self-consistent framework that provides an efficient route for evaluating parameters in metric theories of gravity through black hole shadow geometry. Applied to the Kerr–BR black hole spacetime Podolsky and Ovcharenko (2025), we obtain constraints on the magnetic-field strengths of M87* and Sgr A*; the latter aligns with independent polarization estimates Michail et al. (2023), and both follow the $1/M$ scaling of the horizon-scale magnetization, demonstrating that the observed horizon-scale magnetization can be naturally accommodated within Einstein’s gravity. Therefore, the framework developed here offers a new geometric route to probe near-horizon phenomena and demonstrates that magnetized black hole configurations remain compatible with astrophysical observations within GR, providing a natural foundation for extending such analyses to more general theories of gravity. Beyond GR, theories with additional dynamical fields or couplings introduce a vastly larger parameter space. The adaptive and inference oriented design of our method, which eliminates the need for exhaustive parameter sweeps, makes it particularly well suited for these high complexity models and for systematic observational tests of their signatures. By linking the enlarged theoretical parameter space to measurable image features, this framework offers new opportunities for testing GR and its possible extensions with upcoming EHT and next-generation VLBI arrays.
@@
Acknowledgments.—This work is supported in part by the National Key Research and Development Program of China (Grant No. 2020YFC2201503), the National Natural Science Foundation of China (Grants No. 12475056, No. 12247101, No. 12205243), Gansu Province’s Top Leading Talent Support Plan, the Natural Science Foundation of Gansu Province (No. 22JR5RA389), and the 111 Project (Grant No. B20063).
Appendix A Details
@@
This Appendix details the technical subtleties and calibration steps underlying the main text. For the mapping between pixel coordinates and physically measurable quantities, we establish a normalization procedure based on the Schwarzschild black hole. For a Schwarzschild black hole, the shadow size is absolute and depends only on its mass, independent of the observer’s location or orientation. After setting M=1M\!=\!1, the shadow radius is fixed at $3\sqrt{3}$. In numerical backward ray-tracing simulations, however, the apparent size of the shadow on the image plane depends on the observer’s setup, such as the inclination, field of view $\alpha_{\rm fov}$, and distance $r_{0}$, rather than on the intrinsic physical scale. We therefore use the Schwarzschild case ($\tilde{a}\!=\!\tilde{B}\!=\!0$) as a reference standard: by comparing the simulated image size of the Kerr-BR black hole with that of the Schwarzschild one under identical conditions (same $M$, $\vartheta_{0}$, $\alpha_{\rm fov}$, and $r_{0}$), we can calibrate the apparent shadow scale on the image plane. To simulate the imaging relationship between Earth and the supermassive black holes M87* and Sgr A*, we take the actual source–observer distance as the parameter $r_{0}$. By appropriately adjusting the field of view $\alpha_{\rm fov}$, the simulated Schwarzschild shadows at different distances $r_{0}$ can be made to appear identical in size, effectively serving as a “standard candle” for image normalization. This calibration allows us to disentangle the coupling between the magnetic field strength $\tilde{B}$ and the observational distance when analyzing the Kerr-BR black hole shadows. To determine the quantitative correspondence between $r_{0}$ and $\alpha_{\rm fov}$, we perform a parameter sweep over $(r_{0},\alpha_{\rm fov})$ in the Schwarzschild case and identify the combinations that yield identical pixel-scale shadow diameters in the simulated data.
@@
Figure 5: Contour map of the shadow-area fraction $A_{\rm sh}$ for the Schwarzschild case. Contours show $(r_{0},\,\pi/\alpha_{\rm fov})$ combinations that produce identical apparent shadow sizes, with the contour level representing the fractional shadow area (0–1 scale).
@@
The results of this parameter sweep are shown in Fig. 5, where the mapping is expressed in terms of the inverse field of view parameter $\pi/\alpha_{\rm fov}$. Because $\alpha_{\rm fov}\in(0,\pi)$, the transformed variable spans $(1,\infty)$, providing a more linear and visually uniform representation of the $r_{0}$$\alpha_{\rm fov}$ relation. Each contour represents a set of $(r_{0},\,\pi/\alpha_{\rm fov})$ pairs that produce simulated Schwarzschild shadows of identical apparent size on the image plane. The contour level denotes the shadow-area fraction $A_{\rm sh}$, defined as the ratio of the shadowed region to the total image-plane area, which varies from 0 to 1. Contour levels exceeding $\pi/4\!\approx\!0.785$ are not physically meaningful, as they correspond to configurations where the shadow extends beyond the image boundaries. As expected, larger observer distances correspond to larger values of $\pi/\alpha_{\rm fov}$, reflecting the need for smaller field of view angles to preserve the same apparent shadow diameter. This monotonic correspondence provides a quantitative calibration of the imaging geometry and allows us to express the mapping in analytical form. To facilitate subsequent applications, we model the shadow-area fraction with the empirical power law
$A_{\rm sh}(r_{0},\alpha_{\rm fov})=8.53\left(\frac{\pi}{r_{0}~\alpha_{\rm fov}}\right)^{2}.$ (8)
@@
This empirical model reproduces the numerical contours for $r_{0}\!>\!100M$ as shown in Fig. 5, providing an analytic calibration for subsequent normalization and scaling.
Figure 6: Variation of the Kerr–BR black hole shadow with $\tilde{a}=0.94$ and $\vartheta_{0}=\frac{\pi}{2}$ under different $r_{0}$ and $\tilde{B}$, at a fixed $A_{\rm sh}=\pi/130$.
In recognition of the important contribution by Guo et al. to the study of black hole shadows in this spacetime Wang et al. (2025), we adopt the same observer configuration as used in their analysis—specifically, the setup shown in Fig. 1 of Ref. Wang et al. (2025), with the observer located at $r_{0}\!=\!300M$ and the field of view fixed at $\alpha_{\rm fov}\!=\!\pi/16$. Under these conditions, the apparent shadow of the Schwarzschild black hole occupies a fractional area of $A_{\rm sh}\!=\!0.02413\!\simeq\!\pi/130$ on the image plane. We take this value as a “standard candle” throughout this Letter to calibrate the shadow size in the Kerr–BR spacetime. In Fig. 6, we present the simulated Kerr–BR black hole shadows at different observer distances $r_{0}$. The field-of-view angle $\alpha_{\rm fov}$ is adjusted according to Eq. (8) so that the reference Schwarzschild shadow (purple circle) maintains the same apparent size in all panels, serving as a “standard candle” for direct comparison of magneti field effects. Although not central to our study, the Kerr–BR image in Fig. 6(b) exactly matches Fig. 1 of Ref. Wang et al. (2025), confirming the consistency of our setup. As seen from the blue dashed guides, Figs. 6(b)–6(d) show nearly identical shadows, with larger distances requiring weaker magnetic fields to preserve the same apparent scale. In Fig. 6(a), however, the $P_{C}$ point shifts outward, yielding a more oblate shadow and an increased axial ratio—an effect that becomes pronounced only for strong magnetic fields. These features, already discussed in previous works Wang et al. (2025); Ali and Ghosh (2025), mainly serve here to validate the consistency of our geometric normalization. This confirmation allows us to confidently apply the calibrated relation in Eq. (8) to constrain the parameters of the Kerr–BR spacetime in the following analysis.
@@
In addition, for an approximate estimation of the observable shadow scale, we employ the metric-based relation between the physical shadow radius and its apparent angular size, expressed as $\theta_{s}=R_{s}\frac{\mathcal{M}}{D_{O}}$ Amarilla and Eiroa (2012). Based on our Schwarzschild-calibrated normalization scheme, the corresponding angular radius can be expressed in microarcseconds as
@@
where $R_{s}$ denotes, as previously discussed, the normalized shadow radius on the image plane. It depends on the parameters $(r_{0},\alpha_{\rm fov},\tilde{a},\tilde{B},\vartheta_{0})$, and thus $\theta_{s}$ is also a function of these quantities.
Figure 7: Iterative resolution-shooting process of the adaptive backward ray-tracing algorithm. The base image (a) is progressively refined around the marked boundary regions, yielding high-resolution locating pixels after successive iterations.
Figure 8: Eight boundary neighborhoods are resolved independently and reinserted into the global image to update the locating pixels.
@@
We now turn to the second technical challenge discussed in the main text, which concerns maintaining both numerical precision and computational efficiency in image generation. As illustrated in Fig. 2, the accuracy of the extracted shadow observables, such as $R_{s}$ and $\mathcal{D}_{s}$, is highly sensitive to the image resolution. However, the computational cost increases almost geometrically with higher resolution, making it impractical to generate the large ensemble of images required for the MCMC analysis. To address this issue, we develop an adaptive backward ray-tracing algorithm that dynamically refines the sampling resolution around the shadow boundary.
@@
As shown in Fig. 1, the information required to determine the shadow observables depends solely on the coordinate accuracy of the characteristic points $P_{i}$ (i=A,B,C,Di\!=\!A,B,C,D), implying that a uniformly high image resolution outside these regions is unnecessary. Inspired by the shooting method commonly used in numerical root-finding, we employ a resolution-shooting iteration that begins with a coarse image and progressively refines only the regions relevant to the shadow boundary. In our procedure, the target is to locate the characteristic points $P_{i}$ (i=A,B,C,Di\!=\!A,B,C,D), but they cannot be identified directly on a discrete image grid. Instead, for each $P_{i}$ we first bracket it by two locating endpoints $P_{i_{1}}$ and $P_{i_{2}}$ lying on opposite sides of the shadow edge. These endpoints serve as the operational handles of the algorithm, marking the pixel regions that require refinement and defining the characteristic point as $P_{i}\!\equiv\!\tfrac{1}{2}\,(P_{i_{1}}+P_{i_{2}})$. We start with a base image of $64\times 64$ pixels, as shown in Fig. 7(a), where the initial locating pixels at positions $P^{(0)}_{i_{1}}$ and $P^{(0)}_{i_{2}}$ are marked in orange to identify the boundary regions for subsequent refinement. The base image is treated as a $64\times 64$ matrix with pixel values 0 (white), 1 (black), and 2 (orange for marked regions). Each orange pixel is expanded to cover a local $(\pm n_{1})$ neighborhood, forming eight $7\times 7$ orange subregions for $n_{1}\!=\!3$. Subsequently, all pixels in the image are expanded into $n_{2}\times n_{2}$ submatrices and flattened to generate a super-resolved image of size $(64n_{2})\times(64n_{2})$, as shown in Fig. 7(b), where $n_{1}\!=\!3$ and $n_{2}\!=\!4$. At this stage, each marked local neighborhood contains $(n_{1}n_{2})\times(n_{1}n_{2})\!=\!28\times 28$ pixels. We then initiate the first iteration, performing backward ray-tracing calculations only within these eight $28\times 28$ local neighborhoods, as illustrated in Fig. 8. The results of the eight refined subregions are inserted back into the original image, producing the updated map in Fig. 7(c). This image accurately reproduces the characteristic point information that would be obtained from a full $256\times 256$ ray-tracing simulation, with the updated locating pixels markers denoted as $P^{(1)}_{i_{1}}$ and $P^{(1)}_{i_{2}}$. The procedure is then repeated for a second iteration, raising the local resolution to $1024\times 1024$ and yielding $P^{(2)}_{i_{1}}$ and $P^{(2)}_{i_{2}}$. Within this framework, the iterative refinement allows us to achieve sufficiently high effective resolution to accurately determine the locating pixels $P^{(N)}_{j}$ $(j\!=\!1,\dots,8)$, where $N$ denotes the iteration number and the set $\{P^{(N)}_{j}\}\!=\!\{P^{(N)}_{A_{1}},\dots,P^{(N)}_{D_{2}}\}$. By appropriately choosing $n_{1}$ and $n_{2}$, the algorithm reproduces the same locating pixels as a full all-pixel simulation while requiring only a fraction of its computational resources.
@@
where $E^{(N)}_{k}\!=\!\{E^{(N)}_{1,x},\dots,E^{(N)}_{8,y}\}$ denotes the set of percentage corrections for all coordinate components in the current iteration, computed from the eight locating pixels $P^{(N)}_{j}$ with coordinates $\{x^{(N)}_{j},y^{(N)}_{j}\}$. Each coordinate is normalized by the image resolution so that $\sigma^{(N)}$ becomes a dimensionless indicator of numerical convergence. It quantifies the stability of the locating pixels across iterations, serving as an internal measure of consistency. As the iteration proceeds, $\sigma^{(N)}$ is expected to decrease toward zero, and convergence is achieved when $\lVert\sigma^{(N)}\rVert<\varepsilon$. If $\sigma^{(N)}>\sigma^{(N-1)}$ and $\bar{E}^{(N)}>\bar{E}^{(N-1)}$, the algorithm automatically adjusts the refinement parameters by decreasing $n_{1}$ and increasing $n_{2}$ to re-search the target region. This error-controlled feedback mechanism makes the method inherently adaptive. The adaptive ray-tracing framework developed here provides both accuracy and efficiency for shadow modeling.

### v2

@@ Abstract
Remarks.—In summary, we have constructed a general and self-consistent framework that provides an efficient route for evaluating parameters in metric theories of gravity through black hole shadow geometry. Applying this framework to the Kerr–BR black hole spacetime Podolsky and Ovcharenko (2025), we obtained constraints on the magnetic-field strengths of M87* and Sgr A*. The result for Sgr A* aligns with independent polarization estimates Michail et al. (2023), demonstrating that the observed horizon-scale magnetization can be naturally accommodated within Einstein’s gravity. Therefore, the framework developed here offers a new geometric approach to probing near-horizon phenomena and demonstrates that magnetized black hole configurations remain compatible with astrophysical observations within GR. This provides a natural foundation for extending such analyses to more general theories of gravity. Beyond GR, theories involving additional dynamical fields or couplings introduce a vastly larger parameter space. The adaptive and inference-oriented design of our method, which eliminates the need for exhaustive parameter sweeps, makes it particularly well suited for these high-complexity models and for systematic observational tests of their signatures. By linking an enlarged theoretical parameter space to measurable image features, this framework opens new opportunities for testing GR and its possible extensions with upcoming EHT and next-generation VLBI arrays.
@@
Acknowledgments.—This work is supported in part by the National Natural Science Foundation of China (Grants No. 12475056, No. 12247101, No. 12205243), Gansu Province’s Top Leading Talent Support Plan, the Natural Science Foundation of Gansu Province (No. 22JR5RA389), and the 111 Project (Grant No. B20063).
Appendix A Schwarzschild calibration
This Appendix provides the Schwarzschild calibration scheme employed in the main analysis. For the mapping between pixel coordinates and physically measurable quantities, we establish a normalization procedure based on the Schwarzschild black hole.
A.1 Geometric normalization
@@
For a Schwarzschild black hole, the shadow size is absolute and depends only on its mass, independent of the observer’s location or orientation. After setting M=1M\!=\!1, the physical shadow radius is fixed at $R_{\rm phys}\!=\!3\sqrt{3}$. In numerical ray-tracing simulations, however, the apparent size of the shadow on the image plane depends on the observer’s setup, such as the field of view $\alpha_{\rm fov}$ and distance $r_{0}$ rather than being determined solely by the intrinsic physical scale. We therefore use the Schwarzschild case ($\tilde{a}\!=\!\tilde{B}\!=\!0$) as a reference standard: by comparing the simulated image size of the Kerr-BR black hole with that of the Schwarzschild one under identical conditions, we can calibrate the apparent shadow scale on the image plane. To simulate the imaging relationship between Earth and the supermassive black holes M87* and Sgr A*, we take the actual source–observer distance as the parameter $r_{0}$. By appropriately adjusting the field of view $\alpha_{\rm fov}$, the simulated Schwarzschild shadows at different distances $r_{0}$ can be made to appear identical in size, effectively serving as a “standard candle” for image normalization. This calibration allows us to disentangle the coupling between the magnetic field strength $\tilde{B}$ and the observational distance when analyzing the Kerr-BR black hole shadows.
Figure 5: Schematic illustration of the shadow-area fraction $A_{\rm sh}$ for the Schwarzschild black hole.
Before establishing the quantitative correspondence between $r_{0}$ and $\alpha_{\rm fov}$, we introduce the Schwarzschild shadow-area fraction $A_{\rm sh}$, defined as the ratio of the shadowed region to the total image-plane area, as illustrated in Fig. 5. The left panel shows a schematic example of a discretized image plane (here a $16\times 16$ grid for illustration), in which black pixels denote the shadow region and orange pixels denote light rays reaching the observer. The right panel shows the corresponding two-dimensional binary matrix, where 1 is assigned to shadow pixels and 0 otherwise. The shadow-area fraction is then computed as the normalized sum of the matrix elements, which converges to the physical area fraction in the limit of sufficiently high-resolution. Now, we perform a parameter sweep over $(r_{0},\alpha_{\rm fov})$ in the Schwarzschild case and identify the combinations that yield identical pixel-scale shadow diameters, corresponding to the same Schwarzschild shadow-area fraction $A_{\rm sh}$, in the simulated data.
@@
Figure 6: Contour map of the Schwarzschild shadow-area fraction $A_{\rm sh}$. Contours show $(r_{0},\,\pi/\alpha_{\rm fov})$ combinations that produce identical apparent shadow sizes, with the contour level representing the fractional shadow area (0–1 scale).
@@
The results of this parameter sweep are shown in Fig. 6, where the mapping is expressed in terms of the inverse field of view parameter $\pi/\alpha_{\rm fov}$. Because $\alpha_{\rm fov}\in(0,\pi)$, the transformed variable spans $(1,\infty)$, providing a more linear and visually uniform representation of the $r_{0}$$\alpha_{\rm fov}$ relation. Each contour represents a set of $(r_{0},\,\pi/\alpha_{\rm fov})$ pairs that produce simulated Schwarzschild shadows of identical apparent size on the image plane. The contour level denotes the Schwarzschild shadow-area fraction $A_{\rm sh}$, which varies from 0 to 1. Contour levels exceeding $\pi/4\!\approx\!0.785$ are not physically meaningful, as they correspond to configurations where the shadow extends beyond the image boundaries. As expected, larger observer distances correspond to larger values of $\pi/\alpha_{\rm fov}$, reflecting the need for smaller field of view angles to preserve the same apparent shadow diameter. This monotonic correspondence provides a quantitative calibration of the imaging geometry and allows us to express the mapping in analytical form. To facilitate subsequent applications, we model the shadow-area fraction with the empirical power law
$A_{\rm sh}(r_{0},\alpha_{\rm fov})=8.482\left(\frac{\pi}{r_{0}~\alpha_{\rm fov}}\right)^{2}.$ (8)
@@
This empirical model reproduces the numerical contours for $r_{0}\!>\!100$ as shown in Fig. 6, providing an analytic calibration for subsequent normalization and scaling. Using the shadow-area fraction $A_{\rm sh}$, we obtain the proportional relation
$A_{\rm sh}:\pi R_{s}^{2}\,=\,\pi\,\big(3\sqrt{3}\big)^{2}:\pi\,R_{\rm phys}^{2},$ (9)
where equating the corresponding area ratios yields the mapping between the Schwarzschild-calibrated physical shadow radius $R_{\rm phys}$ and the image-plane radius $R_{s}$.
A.2 Conversion to celestial angular radius
@@
With the Schwarzschild-calibrated physical shadow radius $R_{\rm phys}$ obtained, the final step is to convert this calibrated shadow radius into the celestial angular radius $\theta_{s}$ observable from Earth. To this end, we employ the standard relation between the physical shadow radius and its apparent angular size, given by $\theta_{s}\!=\!R_{\rm phys}\frac{\mathcal{M}}{D_{O}}$ Amarilla and Eiroa (2012). Based on our Schwarzschild-calibrated normalization scheme, the corresponding angular radius can be written in microarcseconds as
@@
where $R_{s}$ denotes, as previously discussed, the radius on the image plane. It depends on the parameters $(r_{0},\alpha_{\rm fov},\tilde{a},\tilde{B},\vartheta_{0})$, and thus $\theta_{s}$ is also a function of these quantities.
A.3 Consistency tests
To assess the robustness of this calibration scheme, we perform consistency tests using the Kerr–BR spacetime as a representative example. In recognition of the important contribution by Guo et al. to the study of black hole shadows in this spacetime Wang et al. (2025), we adopt the same observer configuration as used in their analysis—specifically, the setup shown in Fig. 1 of Ref. Wang et al. (2025), with the observer located at $r_{0}\!=\!300$ and the field of view fixed at $\alpha_{\rm fov}\!=\!\pi/16$.
Under these conditions, the apparent shadow of the Schwarzschild black hole occupies a fractional area of $A_{\rm sh}\!=\!0.02413\!\simeq\!\pi/130$ on the image plane. We take this value as a “standard candle” to calibrate the shadow size in the Kerr–BR spacetime.
Figure 7: Variation of the Kerr–BR black hole shadow with $\tilde{a}=0.94$ and $\vartheta_{0}=\frac{\pi}{2}$ under different values of $r_{0}$ and $\tilde{B}$, at a fixed $A_{\rm sh}=\pi/130$.
In Fig. 7, we present the simulated Kerr–BR black hole shadows at different observer distances $r_{0}$. The field-of-view angle $\alpha_{\rm fov}$ is adjusted according to Eq. (8) so that the reference Schwarzschild shadow (purple circle) maintains the same apparent size in all panels, serving as a “standard candle” for direct comparison of magneti field effects. Although not central to our study, the Kerr–BR image in Fig. 7(b) exactly matches Fig. 1 of Ref. Wang et al. (2025), confirming the consistency of our setup. As indicated by the blue dashed guides, Figs. 7(b)–7(d) exhibit nearly identical shadows, with larger observer distances requiring weaker magnetic fields to reproduce the same apparent scale. This demonstrates a distance–field-strength coupling that our calibration scheme removes. This confirmation allows us to confidently apply the calibrated relation in Eq. (8) to constrain the parameters of the Kerr–BR spacetime in the following analysis.
Appendix B Adaptive ray-tracing algorithm
@@
This Appendix provides the details of the adaptive ray-tracing algorithm used to construct high-resolution black hole shadow images with controlled numerical accuracy. As illustrated in Fig. 2, the accuracy of the extracted shadow observables, such as $R_{s}$ and $\mathcal{D}_{s}$, is highly sensitive to the image resolution. However, the computational cost increases almost geometrically with higher resolution, making it impractical to generate the large ensemble of images required for the MCMC analysis. To address this issue, we develop an adaptive backward ray-tracing algorithm that dynamically refines the sampling resolution around the shadow boundary.
Figure 8: Iterative “resolution-shooting” process of the adaptive backward ray-tracing algorithm. The base image (a) is progressively refined around the marked boundary regions, yielding high-resolution locating pixels after successive iterations.
Figure 9: The eight boundary patches surrounding the shadow edge are refined independently. Each refined subregion is then inserted back into the global image, providing updated locating pixels for the next iteration.
@@
As shown in Fig. 1, the information required to determine the shadow observables depends solely on the coordinate accuracy of the characteristic points $P_{i}$, implying that a uniformly high image resolution outside these regions is unnecessary. Inspired by adaptive mesh refinement techniques used in numerical relativity, we employ a coarse-to-fine “resolution-shooting” iteration that begins with a coarse image and progressively refines only the regions relevant to the shadow boundary. In our procedure, the target is to locate the characteristic points $P_{i}$, but they cannot be identified directly on a discrete image grid. Instead, for each $P_{i}$ we first bracket it by two locating endpoints $P_{i_{1}}$ and $P_{i_{2}}$ lying on opposite sides of the shadow edge. These endpoints serve as the operational handles of the algorithm, marking the pixel regions that require refinement and defining the characteristic point as $P_{i}\!\equiv\!\tfrac{1}{2}\,(P_{i_{1}}+P_{i_{2}})$. This midpoint-based definition ensures that the characteristic points systematically converge toward the true boundary location as the resolution increases. We start with a base image of $64\times 64$ pixels, as shown in Fig. 8(a), where the initial locating pixels at positions $P^{(0)}_{i_{1}}$ and $P^{(0)}_{i_{2}}$ are marked in orange to identify the boundary regions for subsequent refinement. The base image is treated as a $64\times 64$ matrix with pixel values 0 (white), 1 (black), and 2 (orange for marked regions). Each orange pixel is expanded to cover a local $(\pm n_{1})$ neighborhood (controlling the size of the coarse boundary patch), forming eight $7\times 7$ orange subregions for $n_{1}\!=\!3$. Subsequently, all pixels in the image are expanded into $n_{2}\times n_{2}$ submatrices (controlling the local refinement factor) and flattened to generate a super-resolved image with an effective resolution of order $(64n_{2})\times(64n_{2})$, as shown in Fig. 8(b), where $n_{2}\!=\!4$. At this stage, each marked local neighborhood contains $(n_{1}n_{2})\times(n_{1}n_{2})\!=\!28\times 28$ pixels. We then initiate the first iteration, performing backward ray-tracing calculations only within these eight $28\times 28$ local neighborhoods, as illustrated in Fig. 9. The results of the eight refined subregions are inserted back into the original image, producing the updated map in Fig. 8(c). This image accurately reproduces the characteristic point information that would be obtained from a full $256\times 256$ ray-tracing simulation, with the updated locating pixels markers denoted as $P^{(1)}_{i_{1}}$ and $P^{(1)}_{i_{2}}$. The procedure is then repeated for a second iteration, raising the local resolution to $1024\times 1024$ and yielding $P^{(2)}_{i_{1}}$ and $P^{(2)}_{i_{2}}$. Within this framework, the iterative refinement allows us to achieve sufficiently high effective resolution to accurately determine the locating pixels $P^{(N)}_{j}$ $(j\!=\!1,\dots,8)$, where $N$ denotes the iteration number and the set $\{P^{(N)}_{j}\}\!=\!\{P^{(N)}_{A_{1}},\dots,P^{(N)}_{D_{2}}\}$. By appropriately choosing $n_{1}$ and $n_{2}$, the algorithm reproduces the same locating pixels as a full all-pixel simulation while requiring only a fraction of its computational resources.
@@
where $E^{(N)}_{k}\!=\!\{E^{(N)}_{1,x},\dots,E^{(N)}_{8,y}\}$ denotes the set of percentage corrections for all coordinate components in the current iteration, computed from the eight locating pixels $P^{(N)}_{j}$ with coordinates $\{x^{(N)}_{j},y^{(N)}_{j}\}$. The coordinate differences of the locating pixels (i.e., the changes in their $x$- and $y$-positions) are expressed as relative percentages with respect to the previous iteration. This allows $\sigma^{(N)}$ to serve as a dimensionless indicator of numerical convergence, quantifying the stability of the locating pixels across iterations and acting as an internal consistency measure. As the iteration proceeds, $\sigma^{(N)}$ is expected to decrease toward zero, with convergence achieved when $\sigma^{(N)}\!<\!\varepsilon$. If $\sigma^{(N)}\!>\!\sigma^{(N-1)}$ and $\bar{E}^{(N)}\!>\!\bar{E}^{(N-1)}$, the algorithm dynamically adjusts by increasing $n_{1}$ and decreasing $n_{2}$. This adjustment widens the coarse boundary patch while reducing the refinement level until a reliable search region is re-established. Such an error-controlled feedback mechanism makes the method inherently adaptive. Ultimately, the adaptive ray-tracing framework developed here provides both accuracy and efficiency for shadow modeling.